Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.
Please and thank you
Sincerely
Alisha

Answer by nerdybill(7384)

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.
.
Standard form of a line:
Ax + By = C where A is a positive integer, and B, and C are integers
.
First, determine slope between the two points using:
m = (y2-y1)/(x2-x1)
m = (4-3)/(4-1)
m = 1/3 (slope)
.
Next, use one of the two points (1,3) and the slope (1/3) and plug it into the "point-slope" form:
y-y1 = m(x-x1)
y-3 = (1/3)(x-1)
multiplying both sides by 3:
3y-9 = x-1
subtracting x from both sides:
-x + 3y-9 = -1
adding 9 to both sides:
-x + 3y = 8 (this is what they're looking for)